At a given instant, the blood pressure in the heart is 1.50 104 Pa. If an artery in the brain is 0.40 m above the heart, what is the pressure in the artery? Ignore any pressure changes due to blood flow. (See Table 11.1 for appropriate constants.)
Pa
Answer #1

P1 = 1.50*10^4 Pa

P2 = ?

h2 - h1 = 0.40 m

Density of blood ρ = 1060 Kg/m^3

According to bernoullis pricniple

            P2 = P1 + ρg(h1 - h2)

                 = 1.50*10^4 + [1060*9.8*(-0.4)]

                 = 1.5*10^4 - 0.415*10^4

                 = 1.085*10^4 Pa